OpenStudy (anonymous):

Write an inequality describing the range of possible values of x. A. 1/2

OpenStudy (anonymous):



OpenStudy (anonymous):

everything inside of the triangles is accepted solution?

OpenStudy (anonymous):

what ?

OpenStudy (anonymous):

"accepted solution" .. what does that mean?

OpenStudy (anonymous):

the valid range is described with the geometry right? does it state which of the areas are part of the solution and which aren't

OpenStudy (anonymous):

right, and no.

OpenStudy (anonymous):

OK I guess everything inside of the triangles is part of solution and the rest isn't

OpenStudy (anonymous):

? its not

OpenStudy (anonymous):

the lines of the triangles being a border

OpenStudy (anonymous):

oohhh ok

OpenStudy (anonymous):

my answer would be D.

OpenStudy (anonymous):

for example we have 2x+3 as dividing line. it means the function is true for anything that's greater than 2x + 3. for example 3 x would be greater at the same x value, or 2x+5

OpenStudy (anonymous):

oh. ok?

OpenStudy (anonymous):

also if you plot a point at the height y=1 at x position=2, that'd be wrong because y is supposed to be 2x and not smaller.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

So my answers not D then

OpenStudy (anonymous):

the lines divide the solution sets. one set is included and the other is not. to check which one is included, pick any point in either one and see if it is a valid point

OpenStudy (anonymous):

what example point lies inside the lower triangle? something where y is greater than 2x+3. so something where when we put x=2, it does not give 2(2)+3 = 4+3=7 instead it gives a higher value like 12

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

that's the first limitation

OpenStudy (anonymous):

so im left with A,B,C as my answers.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

its probably something with 2 and 3 as well because that's in the equation for the dividing line

OpenStudy (anonymous):

of course we should do it properly then we can easily see which one of them...

OpenStudy (anonymous):

B should be it then because it has 2 & 3.

OpenStudy (anonymous):

C** not B.

OpenStudy (anonymous):

we have already given the y value and have to find x, so that it adds up and in order for it to add up it must have a reverse relationship

OpenStudy (anonymous):

has 2&3.

OpenStudy (anonymous):

y = 2x + 3. so, what does x need to be - re-arranging: y - 3 = 2x x = (y-3)/2

OpenStudy (anonymous):

C is most likely

OpenStudy (anonymous):

thanks! :)

OpenStudy (anonymous):